# Generation of Drug-Induced Cardiac Reactions towards Virtual Clinical Trials

### v1

@@ Appendix A Details of Datasets
Appendix B Additional Visualization Results
Figures 7 presents additional visual experimental results of drug reactions, from which we can observe that after taking Quinidine, the subject’s $T_{peak}-T_{end}$ interval increases from $79\mathrm{ms}$ pre-dose to $142\mathrm{ms}$ post-dose. In our generated ECGs, the $T_{peak}-T_{end}$ interval is $133\mathrm{ms}$. Both values fall within the abnormal range, demonstrating that our method can effectively simulate the reaction of drugs on ECGs.
Figure 8 presents additional visual experimental results with and without EPK and drug reactions, from which we can observe that the introduction of EPK makes the generation of the ECG signal more realistic in detail. In the absence of EPK, the generated ECG signal contains a lot of noise.
Figure 7: Visualization of Quinidine’s reactions on ECGs. Given that lead II is the most suitable for the detection of $T_{peak}-T_{end}$ interval, we visualize a heartbeat from the 10-second ECGs on lead II.

### v2

@@ Appendix B Details of Datasets
Figure 8: Visualization of Quinidine’s reactions on ECGs of lead II.
@@
Appendix C Additional Visualization Results
Figure 8 presents additional visual experimental results of drug reactions, from which we can observe that after taking Quinidine, the subject’s $T_{peak}-T_{end}$ interval increases from $79\mathrm{ms}$ pre-dose to $142\mathrm{ms}$ post-dose. In our generated ECGs, the $T_{peak}-T_{end}$ interval is $153\mathrm{ms}$. Both values fall within the abnormal range, demonstrating that our method can effectively simulate the reaction of drugs on ECGs. Figure 9 presents additional visual experimental results with and without EPK and drug reactions, from which we can observe that the introduction of EPK makes the generation of the ECG signal more realistic in detail. In the absence of EPK, the generated ECG signal contains a lot of noise.